OpenStudy (anonymous):

what is the closest point of x^4+y^4=1 to x+8y=24? (similar to one of the other question that has been written here)

OpenStudy (anonymous):

x= (24-8y) now plug in x make the function 1 a function of x

OpenStudy (anonymous):

differentiate it to find vals of x get the points and find the distances

OpenStudy (anonymous):

sry mistake

OpenStudy (anonymous):

\[\sqrt[4]{1-y ^{4}} - (24-8y)\]

OpenStudy (anonymous):

no (24-8y)^4 + y^4 = 1

OpenStudy (anonymous):

aha u wanna find the point that is for on both of these equation?

OpenStudy (anonymous):

that is on both of these equation*

OpenStudy (anonymous):

Differentiate the fourth-degree equation implicitly and solve for the derivative to get \[y'=-x ^{3}/y ^{3}.\] This is the slope of the line tangent to the curve at (x,y).. The slope of the normal line to the curve there is \[y ^{3}/x ^{3}.\] The line has slope -1/8, and any line perpendicular to it has slope 8, the same slope as the normal line to the curve. So we have\[y ^{3}/x ^{3}=8,\] simplifying to y= 2x. Then \[x ^{4}+(2x)^{4}=1\rightarrow x= 1/\sqrt[4]{17}.\] So \[y = 2/\sqrt[4]{17},\]and the point on the curve nearest the line is \[(1/\sqrt[4]{17},2/\sqrt[4]{17}).\]

OpenStudy (anonymous):

thanks very heplful now i can't understand the way that u found y′=−x3/y3.

OpenStudy (anonymous):

\[4x ^{3} + 4y ^{3}y' = 0\] after implicit differentiation is done. Solve for y'.

OpenStudy (anonymous):

I see thanks abtrehearn

OpenStudy (anonymous):

:^)